The complement of an angle is 6° less than twice the measure of the angle itself. Find the angle and its complement.
Answer by checkley79(3341)

X+Y=180
X=2Y-6
(2Y-6)+Y=180
2Y-6+Y=180
3Y=180+6
3Y=186
Y=186/3
Y=62 ANS. FOR THE SMALLER ANGLE.
X=2*62-6
X=124-6
X=118 ANS FOR THE LARGER ANGLE.
PROOF:
62+118=180
180=180